OpenStudy (anonymous):

how do i do 25 - 8t over 5.

OpenStudy (anonymous):

(25-8t)/5 = 25/5 - 8t/5 = 5 - 8t/5

OpenStudy (anonymous):

or if you meant 25 - 8t/5, then... 25*5/5 - 8t/5 = (125-8t)/5

OpenStudy (anonymous):

thak yoy thanks i love you